OpenStudy (ilovecomputers1):

How is an exponential function created? How can the properties of exponents help solve exponential equations? What are some real-world applications for exponential functions?

OpenStudy (larseighner):

What you need here is a definition. A exponential function is a function in which a variable appears in an exponent. It is NOT when a variable is raised to an exponent. Not exponential: \(\large f(x)=x^\pi,\;f(y)= \sqrt y\) Exponential: \(\large f(x)=\pi^x,\;f(y)= (\sqrt y)^{2y})\) This is an important distinction. For the rest you should do your own thinking.